OpenStudy (anonymous):

which of the following numbers is an example of an irrational number a 1/32 b 3.141592... c √16 or d -22

OpenStudy (anonymous):

Irrational Number: A real number that cannot be written as a simple fraction - the decimal goes on forever without repeating. Pi is an irrational number: 3.141592

OpenStudy (anonymous):

The only number that you can't represent by division of 2 integers \[ \frac{1}{32} \to \text{is already a division of 2 integers, 1 and 32} \\ \sqrt{16} = 4 = \frac{4}{1} \to \text{is already a division of 2 integers, 4 and 1} \\ -22 = \frac{-22}{1} \to \text{is already a division of 2 integers, -22 and 1} \] So we're left with 3.141592 which seems to be like \( \pi \) which is a very known irrational: http://en.wikipedia.org/wiki/Pi

OpenStudy (anonymous):

thank you guys